find the angle between the two st lines 3x=4y+7 and 5y=12x+6 and also the equations to the two lines which pass through the point (4,5) and make equal angles with the two given lines

Grade:12th pass

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below,
4y = 3x - 7........(1)
5y = 12x + 6........(2)
m_{1} = \frac{3}{4}, m_{2} = \frac{12}{5}
Angle between line 1 & 2
tan\theta = |\frac{m_{1}-m_{2}}{1+m_{1}m_{2}}|
tan\theta = |\frac{\frac{3}{4}-\frac{12}{5}}{1+\frac{3}{4}.\frac{12}{5}}|
tan\theta = |\frac{15-48}{20+36}|
tan\theta = \frac{33}{56}
\theta = tan^{-1}\frac{33}{56}
Let the slope of the line be ‘m’
\frac{m-\frac{3}{4}}{1+m.\frac{3}{4}} =\pm \frac{m-\frac{12}{5}}{1+m.\frac{12}{5}}
(4m-3)(12m+5) = \pm (3m+4)(5m-12)
(48m^{2}-16m-15)= \pm (15m^{2}-16m-48)
+ sign is not possible.
48m^{2}-16m-15= -15m^{2}+16m+48
63m^{2}-32m-63= 0
m = \frac{-7}{3}, \frac{3}{7}
Line equations:
(y-5) = \frac{-7}{3}(x-4)
(y-5) = \frac{3}{7}(x-4)
Jitender Singh
9 years ago
Hello Student,
There is a small mistake in the above answer.
Sorry about that
m = -7/9, 9/7